### v1

@@ 3.1 The Invertible Coupling Flow Network
Figure 1: Schematic of the fast-slow neural network (FSNN). The original coordinates are projected into fast-slow coordinates using the invertible neural network (INN). In these coordinates, numerical integration is used to advance the solution to later time steps. The inverse function of the INN is used to convert the fast-slow coordinates into the original coordinate system.
Now we will describe further details of the FSNN architecture. The first component is the invertible function, $h(z;\epsilon)$, which represents the transformation into fast-slow coordinates. The network is based on the invertible coupling flow network developed in [1], where the authors closely follow the architecture provided by [18]. These networks have been shown to be universal approximators to invertible functions in [58]. In this work, we modify the original network by introducing a regularization property which controls the Lipschitz constant of both the forward and inverse map. Such a regularization improves the robustness of the network during training. The constant is a tunable parameter of the network that, similar to the hidden dimensions of the network, can be made arbitrarily large to increase the expressiveness of the network. Therefore, this modification is compatible with the universal approximation result derived in [58]. As such, we introduce the key ingredient of Lipschitz control.
@@
$f(x)=U\Sigma V^{\rm T}x+b,$ (16)
where $U\in\mathbb{R}^{M\times r}$, $V\in\mathbb{R}^{r\times N}$, $\Sigma\in\mathbb{R}^{r\times r}$, $b\in\mathbb{R}^{M}$, $r=\min(M,N)$, $U$ and $V$ are orthogonal matrices satisfying $U^{\rm T}U=V^{\rm T}V=I$, and
$\Sigma={\rm diag}\left[\begin{array}[]{ccc}\sigma_{1}&\dots&\sigma_{r}\end{array}\right],$ (18)
where $\frac{1}{L}\leq\sigma_{i}\leq L$ for $i=1,\dots,r$ and some $L\geq 1$. When $M=N$, the mapping is invertible.
$f^{-1}(x)=V\Sigma^{-1}U^{\rm T}(x-b).$ (19)
@@
The second component of the invertible coupling flow network is the affine coupling flow layer, defined below. We introduce a similar regularization idea to the above.
@@
The affine coupling flow layer is defined by the transformation $y=f(x)$, $x,y\in\mathbb{R}^{M}$, which is implicitly defined by
$\left[\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right]=\left[\begin{array}[]{cc}{\rm diag}(F(x_{2}))&\\ &{\rm diag}(G(y_{1}))\end{array}\right]\left[\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right]+\left[\begin{array}[]{c}B(x_{2})\\ C(y_{1})\end{array}\right],$ (28)
@@
$\left[\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right]=\left[\begin{array}[]{cc}{\rm diag}(F(x_{2}))^{-1}&\\ &{\rm diag}(G(y_{1}))^{-1}\end{array}\right]\left(\left[\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right]-\left[\begin{array}[]{c}B(x_{2})\\ C(y_{1})\end{array}\right]\right).$ (37)
@@
$h=g_{K}\circ f_{K}\circ\dots\circ g_{1}\circ f_{1}$ (38)
@@
$h^{-1}=f_{1}^{-1}\circ g_{1}^{-1}\circ\dots\circ f_{K}^{-1}\circ g_{K}^{-1}.$ (39)
@@
In Theorem 2, the invertible transformation is $\epsilon$-dependent. Therefore, we bake this dependence into the network by including $\epsilon$ as a functional dependence in $F,G,B,C$.
@@ 3.2 The Negative Schur Form Network
$T=\left[\begin{array}[]{cccc}T_{11}&T_{12}&\dots&T_{1p}\\ &T_{22}&\dots&T_{2p}\\ &&\ddots&\vdots\\ &&&T_{pp}\end{array}\right].$ (44)
@@
$\mathcal{M}(R,r,\theta,\phi)=R\left[\begin{array}[]{cc}\cos(\theta)&\sin(\theta)\\ -\sin(\theta)&\cos(\theta)\end{array}\right]+r\left[\begin{array}[]{cc}\cos(\phi)&\sin(\phi)\\ \sin(\phi)&-\cos(\phi)\end{array}\right],$ (49)
@@
$|R|\geq|r|,\quad R<0,\quad\theta\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right).$ (50)
@@
${\rm Re}(\lambda_{\pm})=m\pm\sqrt{\max(m^{2}-p,0)}.$ (52)
Figure 2 shows the regions of the $m{\rm-}p$ plane where both eigenvalues are less than zero (green region), greater than zero (red region), complex (blue region), and real (outside of the blue region). In order for ${\rm Re}(\lambda_{\pm})<0$, we must have $m<0$ and $p>0$. This translates to the constraints in (50).
Figure 2: Plot showing the eigenvalues encountered in various regions of the $m{\rm-}p$ plane, where $m=\frac{1}{2}{\rm Tr}(\mathcal{M})=R\cos(\theta)$ and $p={\rm det}(\mathcal{M})=R^{2}-r^{2}$. The green region corresponds to linear stability, ${\rm Re}(\lambda_{+},\lambda_{-})<0$, and the red region corresponds to linear instability, ${\rm Re}(\lambda_{+},\lambda_{-})>0$. Above the parabola $p=m^{2}$ (in the blue region), eigenvalues are complex. Below the parabola (outside of the blue region), the eigenvalues are real.
@@
$R=-|\bar{r}|e^{|\bar{R}|},\quad r=\bar{r},\quad\theta=\frac{\pi}{2}\tanh(\bar{\theta}),\quad\phi=\bar{\phi}.$ (53)
@@
3.3 The Low-Rank Bilinear Form Network
Finally, we introduce an approximation for the bilinear form term, $B$.
@@ 3.3 The Low-Rank Bilinear Form Network
The low-rank bilinear form network is a universal approximator to the set of bilinear forms.
Consider the bilinear form, $\tilde{B}:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{y}}\rightarrow\mathbb{R}^{N_{z}}$. We can represent the operator of this function as $\tilde{B}_{ijk}$ where $B_{i}\in\mathbb{R}^{N_{x}\times N_{y}}$ for $i=1,\dots,N_{z}$. Let $p=\min(N_{x},N_{y})$, the we can use the singular value decomposition to represent the matrix $\tilde{B}_{i}$ as
$\tilde{B}_{i}=\sum_{r=1}^{p}\sigma_{i}^{(r)}u_{i}^{(r)}(v_{i}^{(r)})^{T},$ (55)
@@
$x^{T}\tilde{B}_{i}y$ $=\sum_{r=1}^{p}\sigma_{i}^{(r)}x^{T}u_{i}^{(r)}(v_{i}^{(r)})^{T}y,=\sum_{r=1}^{p}\sigma_{i}^{(r)}\left(\sum_{j=1}^{N_{x}}u_{ij}^{(r)}x_{j}\right)\left(\sum_{k=1}^{N_{y}}v_{ik}^{(r)}y_{k}\right),$ (56)
@@
$x^{T}\tilde{B}_{i}y$ $=\sum_{r=1}^{p}\left(\sum_{j=1}^{N_{x}}C_{ij}^{(r)}x_{j}\right)\left(\sum_{k=1}^{N_{y}}D_{ik}^{(r)}y_{k}\right).$ (57)
@@
$\tilde{B}(x,y)$ $=\sum_{r=1}^{p}(C^{(r)}x)\odot(D^{(r)}y).$ (58)
The network uses a truncated version of this sum to approximate the action of the bilinear operator. The approximation error of the low-rank bilinear form network is $\|B_{i}-\tilde{B}_{i}\|_{2}=\sigma_{i}^{(R+1)}$. This error can be decreased by increasing $R$.
@@ 3.4 Numerical Integration
The Schur form, $T$ will have eigenvalues that can be of arbitrarily large magnitude. Due to this, the dynamics of (2) will be stiff. Therefore, an Implicit-Explicit (IMEX) time integrator can be used to solve (2) efficiently. The IMEX scheme has been applied to a neural ODE framework in [62]. As shown below, an additional benefit of the proposed FSNN is to eliminiate the need for iterative or direct linear solvers, thereby simplifying the training process.
Consider an IMEX Runge-Kutta method consisting of a diagonally implicit for treating the stiff linear term and an explicit method for treating the nonlinear term. The method can be represented by the following Butcher tableau.
@@
$X^{i}$ $=x_{n-1}+\epsilon\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{g(X^{j},Y^{j},\epsilon)},$ (59a)
$Y^{i}$ $=y_{n-1}+\Delta t\sum_{j=1}^{i}a_{ij}{T(X^{j})Y^{j}}+\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{f(X^{j},Y^{j},\epsilon)}$ (59b)
for $i=1,\dots,s$. The above equation can be written to expose the linear system that needs to be solve.
$Y^{i}$ $=(I-\Delta ta_{ii}T(X^{i}))^{-1}\left(y_{n-1}+\Delta t\sum_{j=1}^{i-1}\left(a_{ij}{T(X^{j})Y^{j}}+\alpha_{ij}{f(X^{j},Y^{j},\epsilon)}\right)\right)$ (60)
@@
$x_{n}$ $=x_{n-1}+\epsilon\Delta t\sum_{i=1}^{s}\beta_{i}{g(X^{i},Y^{i},\epsilon)},$ (61a)
$y_{n}$ $=y_{n-1}+\Delta t\sum_{i=1}^{s}(b_{i}{T(X^{i})Y^{i}}+\beta_{i}{f(X^{i},Y^{i},\epsilon)}).$ (61b)
@@
$x_{n}$ $=x_{n-1}+\epsilon\Delta t\sum_{i=1}^{s}\beta_{i}{g(X^{i},0,\epsilon)},$ (62a)
$X^{i}$ $=x_{n-1}+\epsilon\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{g(X^{j},0,\epsilon)},\qquad i=1,\dots,s.$ (62b)
In this case, one can take larger time steps (e.g. $\Delta t\sim\frac{1}{\epsilon}$) and step over the fast dynamics formally. A hybrid scheme can also be used where the full system in (3.4) can be integrated until $\|y\|$ is sufficiently small, after which slow integration in (3.4) is pursued using larger time steps. We note that some of the first works to use Runge-Kutta based residual networks to learn equations from time series was pursued in [51, 50].
@@
A closure discovery procedure is implied through the above integration with large time steps. For a given full state vector $z_{0}$, the transformation $h$ first maps to the separate state of $(x_{0},y_{0})$. The state vector is then projected onto the closure of $y=0$; the system for $x$ is thus closed and evolved using (3.4). Finally, the evolved state of $x$ is lifted back to the full state $z$ using $h^{-1}$.
@@ 4 Numerical Implementation
Figure 3: Schematic of training the fast-slow neural network (FSNN). The network is trained on dynamics data, $\mathcal{D}$, which contains examples of trajectories on both the fast and slow timescales and manifold data, $\mathcal{M}$, which contains examples of solution vectors that lie on the slow manifold. The overall loss is a combination of the dynamics loss, which includes $L_{2}$ losses in the original coordinates ($\mathcal{L}_{\rm system}$), fast coordinates ($\mathcal{L}_{\rm fast}$), and slow coordinates ($\mathcal{L}_{\rm slow}$), and the manifold loss ($\mathcal{L}_{\rm manifold}$), which is an $L_{2}$ loss of the fast coordinate, which should map to $y=0$ on the slow manifold.
In this section, we describe the numerical implementation of the fast-slow neural network and its corresponding training algorithm. The computation of the loss function and its gradient is summarized in Figure 3. The training set consists of both trajectory data, which describes the behavior of the dynamics, and manifold data, which describes the limiting behavior of the system. The trajectory dataset, $\mathcal{D}$, contains time series, $\{(t_{n}^{(k)},z_{n}^{(k)})\}_{n=1}^{N_{t}}$, and parameters $\epsilon^{(k)}$, for $k=1,\dots,N_{d}$, where $N_{d}$ is the number of trajectory data samples. The manifold dataset, $\mathcal{M}$, contains multiple examples of the solution on the slow manifold, $z^{(k)}$, with corresponding parameters $\epsilon^{(k)}$, for $k=1,\dots,N_{m}$. Examples of manifold data can be obtained either experimentally by integrating a model for a sufficiently long time with a small $\epsilon$ or analytically if the equations permit an $\epsilon=0$ reduction or asymptotic expansion.
@@
$\mathcal{L}=\mathcal{L}_{\rm system}+\mathcal{L}_{\rm fast}+\mathcal{L}_{\rm slow}+\mathcal{L}_{\rm manifold}.$ (63)
@@
$L(t,z,\tilde{z})=\sum_{n=0}^{N_{t}-1}\frac{1}{2}\left((z_{n+1}-\tilde{z}_{n+1})^{2}+(z_{n}-\tilde{z}_{n})^{2}\right)(t_{n+1}-t_{n}).$ (64)
@@
$\mathcal{L}_{\rm system}$ $=\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}L(t,z^{(k)},\tilde{z}^{(k)}),$ (65a)
$\mathcal{L}_{\rm fast}$ $=\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}L(t,y^{(k)},\tilde{y}^{(k)}),$ (65b)
$\mathcal{L}_{\rm slow}$ $=\frac{1}{\epsilon^{(k)}N_{d}}\sum_{k=1}^{N_{d}}L(t,x^{(k)},\tilde{x}^{(k)}).$ (65c)
@@
In practice, we split the trajectory dataset, $\mathcal{D}$, into a fast-scale subset, with time steps chosen to smoothly capture the fast dynamics converging to the slow manifold, and a slow-scale subset, with larger time steps chosen to capture the variation of the slow variables after the dynamics sufficiently close to the slow manifold. The trajectories in fast-scale subset are treated using the full integration in (3.4) while the trajectories from the slow-scale subset are treated using the slow-manifold integration in (3.4).
@@
$\mathcal{L}_{\rm manifold}=\frac{1}{N_{m}}\sum_{k=1}^{N_{m}}(y^{(k)})^{2}$ (66)

### v2

@@ 3.1 The Invertible Coupling Flow Network
Figure 1: Schematic of the fast-slow neural network (FSNN). The original coordinates are projected into fast-slow coordinates using the invertible coupling flow network. In these coordinates, numerical integration of the Fenichel normal form (8) is used to advance the solution to later time steps. The inverse function of the invertible coupling flow network is used to convert the trajectories in the fast-slow coordinates into the original coordinate system.
Now we will describe further details of the FSNN architecture. The first component is the invertible function, $h(z;\epsilon)$, which represents the transformation into fast-slow coordinates. The network is based on the invertible coupling flow network developed in [54], where the authors closely follow the architecture provided by [55]. This network is an example of an invertible neural network (INN). In particular, invertible coupling flow networks have been shown to be universal approximators to invertible functions in [37]. In this work, we modify the original network by introducing a regularization property which controls the Lipschitz constant of both the forward and inverse map. Such a regularization improves the robustness of the network during training. The constant is a tunable parameter of the network that, similar to the hidden dimensions of the network, can be made arbitrarily large to increase the expressiveness of the network. Therefore, this modification is compatible with the universal approximation result derived in [37]. As such, we introduce the key ingredient of Lipschitz control.
@@
(15) $f(x)=U\Sigma V^{\rm T}x+b,$
where $U\in\mathbb{R}^{M\times r}$, $V\in\mathbb{R}^{N\times r}$, $\Sigma\in\mathbb{R}^{r\times r}$, $b\in\mathbb{R}^{M}$, $r=\min(M,N)$, $U$ and $V$ are orthogonal matrices satisfying $U^{\rm T}U=V^{\rm T}V=I$, and
(17) $\Sigma={\rm diag}\left[\begin{array}[]{ccc}\sigma_{1}&\dots&\sigma_{r}\end{array}\right],$
where $\frac{1}{L}\leq\sigma_{i}\leq L$ for $i=1,\dots,r$ and some $L\geq 1$.
We remark that when $M=N$, the mapping is invertible.
(18) $f^{-1}(x)=V\Sigma^{-1}U^{\rm T}(x-b).$
@@
The second component of the invertible coupling flow network is the affine coupling flow layer. We adapt the definition provided in Section 2.1 of [37] and introduce a similar regularization idea to the above.
@@
The affine coupling flow layer is defined by the transformation $y={{g}}(x)$, $x,y\in\mathbb{R}^{M}$, which is implicitly defined by
(27) $\left[\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right]=\left[\begin{array}[]{cc}{\rm diag}(F(x_{2}))&\\ &{\rm diag}(G(y_{1}))\end{array}\right]\left[\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right]+\left[\begin{array}[]{c}B(x_{2})\\ C(y_{1})\end{array}\right],$
@@
(36) $\left[\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right]=\left[\begin{array}[]{cc}{\rm diag}(F(x_{2}))^{-1}&\\ &{\rm diag}(G(y_{1}))^{-1}\end{array}\right]\left(\left[\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right]-\left[\begin{array}[]{c}B(x_{2})\\ C(y_{1})\end{array}\right]\right).$
@@
(37) $h=g_{K}\circ f_{K}\circ\dots\circ g_{1}\circ f_{1}$
@@
(38) $h^{-1}=f_{1}^{-1}\circ g_{1}^{-1}\circ\dots\circ f_{K}^{-1}\circ g_{K}^{-1}.$
@@
In Theorem 8, the invertible transformation is $\epsilon$-dependent. Therefore, we incorporate this dependence into the network by including $\epsilon$ as a functional dependence in $F,G,B,C$.
@@ 3.2 The Negative Schur Form Network
(43) $T=\left[\begin{array}[]{cccc}T_{11}&T_{12}&\dots&T_{1p}\\ &T_{22}&\dots&T_{2p}\\ &&\ddots&\vdots\\ &&&T_{pp}\end{array}\right].$
@@
(48) $\mathcal{M}(R,r,\theta,\phi)=R\left[\begin{array}[]{cc}\cos(\theta)&\sin(\theta)\\ -\sin(\theta)&\cos(\theta)\end{array}\right]+r\left[\begin{array}[]{cc}\cos(\phi)&\sin(\phi)\\ \sin(\phi)&-\cos(\phi)\end{array}\right],$
@@
(49) $|R|\geq|r|,\quad R<0,\quad\theta\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right),\quad{{\phi\in\mathbb{R}}}$
@@
(51) ${\rm Re}(\lambda_{\pm})=m\pm\sqrt{\max(m^{2}-p,0)}.$
Figure 2 shows the regions of the $m{\rm-}p$ plane where both eigenvalues are less than zero (green region), greater than zero (red region), complex (blue region), and real (outside of the blue region). In order for ${\rm Re}(\lambda_{\pm})<0$, we must have $m<0$ and $p>0$. In order to enforce these constraints, we enforce the restrictions in (49). We remark that for the first constraint, $R\cos(\theta)<0$, $R$ and $\cos(\theta)$ must have opposite signs. If we focus on the region $\theta\in(-\pi/2,\pi/2)$ where $\cos(\theta)>0$, we would require $R<0$. If the region of $\theta$ were instead shifted by $\pi$, then we would require $R<0$. However, since $\mathcal{M}(R,r,\theta+\pi,\phi)=\mathcal{M}(-R,r,\theta,\phi)$, these choices are equivalent.
Figure 2: Plot showing the eigenvalues encountered in various regions of the $m$-$p$ plane, where $m=\frac{1}{2}{\rm Tr}(\mathcal{M})=R\cos(\theta)$ and $p={\rm det}(\mathcal{M})=R^{2}-r^{2}$. The green region corresponds to linear stability, ${\rm Re}(\lambda_{+},\lambda_{-})<0$, and the red region corresponds to linear instability, ${\rm Re}(\lambda_{+},\lambda_{-})>0$. Above the parabola $p=m^{2}$ (in the blue region), eigenvalues are complex. Below the parabola (outside of the blue region), the eigenvalues are real.
@@
(52) $R=-|\bar{r}|e^{|\bar{R}|},\quad r=\bar{r},\quad\theta=\frac{\pi}{2}\tanh(\bar{\theta}),\quad\phi=\bar{\phi}.$
@@
3.3 The Low-Rank bilinear map Network
Finally, we introduce an approximation for the bilinear map term, $B$.
@@ 3.3 The Low-Rank bilinear map Network
The low-rank bilinear map network is a universal approximator to the set of bilinear forms.
Consider the bilinear form, $\tilde{B}:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{y}}\rightarrow\mathbb{R}^{N_{z}}$. The tensor representation of $\tilde{B}$ is written as $\tilde{B}_{ijk}$ where $\tilde{B}_{i}\in\mathbb{R}^{N_{x}\times N_{y}}$ for $i=1,\dots,N_{z}$. Let $p=\min(N_{x},N_{y})$, then we can use the singular value decomposition to represent the matrix $\tilde{B}_{i}$ as
(54) $\tilde{B}_{i}=\sum_{r=1}^{p}\sigma_{i}^{(r)}u_{i}^{(r)}(v_{i}^{(r)})^{T},$
@@
(55) $x^{T}\tilde{B}_{i}y$ $=\sum_{r=1}^{p}\sigma_{i}^{(r)}x^{T}u_{i}^{(r)}(v_{i}^{(r)})^{T}y,=\sum_{r=1}^{p}\sigma_{i}^{(r)}\left(\sum_{j=1}^{N_{x}}u_{ij}^{(r)}x_{j}\right)\left(\sum_{k=1}^{N_{y}}v_{ik}^{(r)}y_{k}\right),$
@@
(56) $x^{T}\tilde{B}_{i}y$ $=\sum_{r=1}^{p}\left(\sum_{j=1}^{N_{x}}C_{ij}^{(r)}x_{j}\right)\left(\sum_{k=1}^{N_{y}}D_{ik}^{(r)}y_{k}\right).$
@@
(57) $\tilde{B}(x,y)$ $=\sum_{r=1}^{p}(C^{(r)}x)\odot(D^{(r)}y).$
The network uses a truncated version of this sum in equation (53) to approximate the action of the bilinear map $\tilde{B}$. The approximation error of the low-rank bilinear map network is $\|B_{i}-\tilde{B}_{i}\|_{2}=\sigma_{i}^{(R+1)}$. This error can be decreased by increasing $R$.
@@ 3.4 Numerical Integration
The Schur form, $T$ will have eigenvalues that can be of arbitrarily large magnitude. Due to this, the dynamics of (8) will be stiff. Therefore, an Implicit-Explicit (IMEX) time integrator can be used to solve (8) efficiently. The IMEX scheme has been applied to a neural ODE framework in [57]. As shown below, an additional benefit of the proposed FSNN is to eliminate the need for iterative or direct linear solvers, thereby simplifying the training process.
Consider an IMEX Runge-Kutta method consisting of a diagonally implicit method for treating the stiff linear term and an explicit method for treating the nonlinear term. The method can be represented by the following Butcher tableau [58]
@@
(58a) $X^{i}$ $=x_{n-1}+\epsilon\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{g(X^{j},Y^{j},\epsilon)},$
(58b) $Y^{i}$ $=y_{n-1}+\Delta t\sum_{j=1}^{i}a_{ij}{T(X^{j})Y^{j}}+\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{f(X^{j},Y^{j},\epsilon)}$
for $i=1,\dots,s$. In our notation, we take the above summation when $i=1$ to be 0 (i.e., $\sum_{j=1}^{0}(\cdot)=0$). Equation (58b) can be rewritten to solve for $Y^{i}$.
(59) $Y^{i}$ $=(I-\Delta ta_{ii}T(X^{i}))^{-1}\left(y_{n-1}+\Delta t\sum_{j=1}^{i-1}\left(a_{ij}{T(X^{j})Y^{j}}+\alpha_{ij}{f(X^{j},Y^{j},\epsilon)}\right)\right)$
@@
(60a) $x_{n}$ $=x_{n-1}+\epsilon\Delta t\sum_{i=1}^{s}\beta_{i}{g(X^{i},Y^{i},\epsilon)},$
(60b) $y_{n}$ $=y_{n-1}+\Delta t\sum_{i=1}^{s}(b_{i}{T(X^{i})Y^{i}}+\beta_{i}{f(X^{i},Y^{i},\epsilon)}).$
@@
(61a) $x_{n}$ $=x_{n-1}+\epsilon\Delta t\sum_{i=1}^{s}\beta_{i}{g(X^{i},0,\epsilon)},$
(61b) $X^{i}$ $=x_{n-1}+\epsilon\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{g(X^{j},0,\epsilon)},\qquad i=1,\dots,s.$
In this case, one can take larger time steps (e.g. $\Delta t\sim\frac{1}{\epsilon}$) and step over the fast dynamics formally. A hybrid scheme can also be used where the full system in (60) can be integrated until $\|y\|$ is sufficiently small, after which slow integration in (61) is pursued using larger time steps. We note that some of the first works to use Runge-Kutta based residual networks to learn equations from time series was pursued in [59, 60].
@@
A closure discovery procedure is implied through the above integration with large time steps. For a given full state vector $z_{0}$, the transformation $h$ first maps to the separate state of $(x_{0},y_{0})$. The state vector is then projected onto the closure of $y=0$; the system for $x$ is thus closed and evolved using (61). Finally, the evolved state of $x$ is lifted back to the full state $z$ using $h^{-1}$.
@@ 4 Numerical Implementation
Figure 3: Schematic of training the fast-slow neural network (FSNN). The network is trained on dynamics data, $\mathcal{D}$, which contains examples of trajectories on both the fast and slow timescales and manifold data, $\mathcal{M}$, which contains examples of solution vectors that lie on the slow manifold. This data interacts both with the full network and the component involving the invertible neural network (INN). The overall loss is a combination of the dynamics loss, which includes $L_{2}$ losses in the original coordinates ($\mathcal{L}_{\rm system}$), fast coordinates ($\mathcal{L}_{\rm fast}$), and slow coordinates ($\mathcal{L}_{\rm slow}$), and the manifold loss ($\mathcal{L}_{\rm manifold}$), which is an $L_{2}$ loss of the fast coordinate, which should map to $y=0$ on the slow manifold.
In this section, we describe the numerical implementation of the fast-slow neural network and its corresponding training algorithm. The computation of the loss function and its gradient is summarized in Figure 3. The training set consists of both trajectory data, which describes the behavior of the dynamics, and manifold data, which describes the limiting behavior of the system. The trajectory dataset, $\mathcal{D}$, contains time series, $\{(t_{n}^{(k)},z_{n}^{(k)})\}_{n=1}^{N_{t}}$, and parameters $\epsilon^{(k)}$, for $k=1,\dots,N_{d}$, where $N_{d}$ is the number of trajectory data samples. The manifold dataset, $\mathcal{M}$, contains multiple examples of the solution on the slow manifold, $z^{(k)}$, with corresponding parameters $\epsilon^{(k)}$, for $k=1,\dots,N_{m}$. Examples of manifold data can be obtained either experimentally by integrating a model for a sufficiently long time with a small $\epsilon$ or analytically if the equations permit an $\epsilon=0$ reduction or asymptotic expansion. For many examples using the zeroth order expansion with $\epsilon=0$ is sufficient. Often obtaining a higher order expansion is nontrivial.
@@
(62) $\mathcal{L}=\mathcal{L}_{\rm system}+\mathcal{L}_{\rm fast}+\mathcal{L}_{\rm slow}+\mathcal{L}_{\rm manifold}.$
@@
(63) $L(t,z,\tilde{z})=\sum_{n=0}^{N_{t}-1}\frac{1}{2}\left((z_{n+1}-\tilde{z}_{n+1})^{2}+(z_{n}-\tilde{z}_{n})^{2}\right)(t_{n+1}-t_{n}).$
@@
(64a) $\mathcal{L}_{\rm system}$ $=\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}L(t,z^{(k)},\tilde{z}^{(k)}),$
(64b) $\mathcal{L}_{\rm fast}$ $=\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}L(t,y^{(k)},\tilde{y}^{(k)}),$
(64c) $\mathcal{L}_{\rm slow}$ $={{\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}\frac{L(t,x^{(k)},\tilde{x}^{(k)})}{\epsilon^{(k)}}.}}$
@@
In practice, we split the trajectory dataset, $\mathcal{D}$, into a fast-scale subset, with time steps chosen to smoothly capture the fast dynamics converging to the slow manifold, and a slow-scale subset, with larger time steps chosen to capture the variation of the slow variables after the dynamics sufficiently close to the slow manifold. The trajectories in fast-scale subset are treated using the full integration in (60) while the trajectories from the slow-scale subset are treated using the slow-manifold integration in (61).
@@
(65) $\mathcal{L}_{\rm manifold}=\frac{1}{N_{m}}\sum_{k=1}^{N_{m}}(y^{(k)})^{2}$